Two oppositely charged but otherwise identical conductingplates of area 2.50 square centimeters are separated by adielectric 1.80 millimeters thick, with a dielectric constant ofK=3.60. The resultant electric field in the dielectric is1.20 \times 10^{6} volts per meter.
a) Compute the magnitude of the charge per unit areasigma on the conducting plate.
b) Compute the magnitude of the charge per unit areasigma_1 on the surfaces of the dielectric.
c) Find the total electric-field energy U stored in the capacitor.
Answer #1
Concepts and reason

Use the concept of Capacitance of a capacitor with dielectric to solve this problem.

The magnitude of the charge per unit area on the conducting plate is calculated using the expression of charge density.

The magnitude of charge per unit area on the surface of dielectric plate is calculated using the charge density of conducting plate.

The total electric - field energy stored in the capacitor is calculated using the expression for energy stored in the capacitor.

Fundamentals

The expression of charge per unit area on the conducting plate is,

σ=kε0E\sigma = k{\varepsilon _0}E

Here, kk is the dielectric constant, EE is the magnitude of electric field, ε0{\varepsilon _0} is the permittivity of free space, and σ\sigma is the charge density.

The expression for magnitude of charge per unit area on the surface of dielectric plate is,

σ=σ(11k)\sigma ' = \sigma \left( {1 - \frac{1}{k}} \right)

Here, σ\sigma is the charge density of conducting plate σ1{\sigma _1} is the charge density on surface of dielectric plate, and k is the dielectric constant.

The expression to calculate the total energy stored in the capacitor is,

U=12kε0E2AdU = \frac{1}{2}k{\varepsilon _0}{E^2}Ad

Here, kk is the dielectric constant, A is the area, d is the thickness of the plates, E is the magnitude of electric field, ε0{\varepsilon _0}is the permittivity of free space, and UUis the energy stored in the capacitor.

(a)

Calculate the magnitude of the charge per unit area on the conducting plate.

The expression of charge per unit area on the conducting plate is,

σ=kε0E\sigma = k{\varepsilon _0}E

Here, kk is the dielectric constant, EE is the magnitude of electric field, ε0{\varepsilon _0} is the permittivity of free space, and σ\sigma is the charge density per unit area.

Substitute 3.60 for k , 8.85×1012C2/Nm28.85 \times {10^{ - 12}}\,{{\rm{C}}^{\rm{2}}}{\rm{/N}} \cdot {{\rm{m}}^{\rm{2}}} for ε0{\varepsilon _0}, and 1.20×106V/m1.20 \times {10^6}\,{\rm{V/m}}for EE.

σ=(3.60)(8.85×1012C2/Nm2)(1.20×106V/m)=3.82×105C2/Nm2\begin{array}{c}\\\sigma = \left( {3.60} \right)\left( {8.85 \times {{10}^{ - 12}}\,{{\rm{C}}^{\rm{2}}}{\rm{/N}} \cdot {{\rm{m}}^{\rm{2}}}} \right)\left( {1.20 \times {{10}^6}\,{\rm{V/m}}} \right)\\\\ = 3.82 \times {10^{ - 5}}\,\,{{\rm{C}}^{\rm{2}}}{\rm{/N}} \cdot {{\rm{m}}^{\rm{2}}}\\\end{array}

(b)

Calculate the magnitude of charge per unit area on the surface of dielectric plate.

The expression for magnitude of charge per unit area on the surface of dielectric plate is,

σ=σ(11k)\sigma ' = \sigma \left( {1 - \frac{1}{k}} \right)

Here, σ\sigma is the charge density of conducting plate σ1{\sigma _1} is the charge density on surface of dielectric plate, and k is the dielectric constant.

Substitute 3.82×105C2/Nm23.82 \times {10^{ - 5}}\,{{\rm{C}}^{\rm{2}}}{\rm{/N}} \cdot {{\rm{m}}^{\rm{2}}} for σ\sigma and 3.603.60 for k.

σ=3.82×105C2/Nm2(113.60)=2.76×105C/m2\begin{array}{c}\\\sigma ' = 3.82 \times {10^{ - 5}}\,{{\rm{C}}^{\rm{2}}}{\rm{/N}} \cdot {{\rm{m}}^{\rm{2}}}\left( {1 - \frac{1}{{3.60}}} \right)\\\\ = 2.76 \times {10^{ - 5}}\,{\rm{C/}}{{\rm{m}}^{\rm{2}}}\\\end{array}

(c)

Calculate the total electric- field energy stored in the capacitor by using the expression as follows:

U=12kε0E2AdU = \frac{1}{2}k{\varepsilon _0}{E^2}Ad

Convert the area from centimeter square to meter square by multiplying it with104{10^{ - 4}}.

A=(2.5cm2)(104m21cm2)=2.5×104m2\begin{array}{c}\\A = \left( {2.5\,\,{\rm{c}}{{\rm{m}}^{\rm{2}}}} \right)\left( {\frac{{{{10}^{ - 4}}\,{{\rm{m}}^{\rm{2}}}}}{{1\,\,{\rm{c}}{{\rm{m}}^{\rm{2}}}}}} \right)\\\\ = 2.5 \times {10^{ - 4}}\,\,{{\rm{m}}^{\rm{2}}}\\\end{array}

Convert the diameter from millimeter to meter by multiplying it with 103{10^{ - 3}}.

d=(1.8mm)(103m1mm)=1.8×103m\begin{array}{c}\\d = \left( {1.8\,{\rm{mm}}} \right)\left( {\frac{{{{10}^{ - 3}}\,{\rm{m}}}}{{1\,\,{\rm{mm}}}}} \right)\\\\ = 1.8 \times {10^{ - 3}}\,{\rm{m}}\\\end{array}

Substitute 3.60 for k , 8.85×1012C2/Nm28.85 \times {10^{ - 12}}\,{{\rm{C}}^{\rm{2}}}{\rm{/N}} \cdot {{\rm{m}}^{\rm{2}}} for ε0{\varepsilon _0},1.20×106V/m1.20 \times {10^6}\,{\rm{V/m}}for EE, 2.5×104m22.5 \times {10^{ - 4}}\,{{\rm{m}}^{\rm{2}}} for AA, and 1.8×103m1.8 \times {10^{ - 3}}\,{\rm{m}} for dd in expression U=12kε0E2AdU = \frac{1}{2}k{\varepsilon _0}{E^2}Ad.

U=12(3.60)(8.85×1012C2/Nm2)(1.20×106V/m)2(2.5×104m2)(1.8×103m)=1.03×105J\begin{array}{c}\\U = \frac{1}{2}\left( {3.60} \right)\left( {8.85 \times {{10}^{ - 12}}\,{{\rm{C}}^{\rm{2}}}{\rm{/N}} \cdot {{\rm{m}}^{\rm{2}}}} \right){\left( {1.20 \times {{10}^6}\,{\rm{V/m}}} \right)^2}\left( {2.5 \times {{10}^{ - 4}}\,{{\rm{m}}^{\rm{2}}}} \right)\left( {1.8 \times {{10}^{ - 3}}\,{\rm{m}}} \right)\\\\ = 1.03 \times {10^{ - 5}}\,{\rm{J}}\\\end{array}

Ans: Part a

The magnitude of charge per unit area on the conducting plate is 3.82×105C2/Nm23.82 \times {10^{ - 5}}\,{{\rm{C}}^{\rm{2}}}{\rm{/N}} \cdot {{\rm{m}}^{\rm{2}}}.

Part b

The magnitude of charge per unit area on the surface of dielectric plate is 2.76×105C/m22.76 \times {10^{ - 5}}\,{\rm{C/}}{{\rm{m}}^{\rm{2}}}.

Part c

The total electric field-energy stored in the capacitor is 1.03×105J1.03 \times {10^{ - 5}}\,{\rm{J}}.